# Let Real Images be as a Judger, Spotting Fake Images Synthesized with Generative Models

### v1

@@ 4.1 Natural Trace Representation Learning
Formulation. We assume access to a large-scale dataset of real images $D_{r}$. Sample $x\in D_{r}$ is a augmented sample from a image $x_{r}\in\mathbb{R}^{H\times W\times 3}$ (of height $H$, and width $W$). As shown in Figure 2(a), our architecture consists of a feature encoder followed by two projection heads that map real image embeddings into homogeneous and heterogeneous features.
Specifically, feature encoder, $E$, producing feature embedding $e=E(x_{r})$ from the inputs, then decouples the $e$ into homogeneous and heterogeneous features, $z^{hom}=f^{hom}(e)\in\mathbb{R}^{C}$ and $z^{het}=f^{het}(e)\in\mathbb{R}^{C}$ through two projection heads, $f^{hom}$ and $f^{het}$, respectively, where $C$ is the dimensionality of the features. Let $i\in I\equiv\{1...2N\}$ be the index of an arbitrary augmented sample, where there are a total of N samples in a batch, each with two random augmentations ($2N$). The self-supervised contrastive loss of heterogeneous feature representations in real images can be formulated as follows:
@@
where the index $i$ is called the anchor, the index $p$ is called the positive, $\tau$ is a temperature hyperparameter and $K(i)\equiv I\backslash\{i\}$, which contains all the augmented samples except $i$ in a batch. Furthermore, the loss of homogeneous feature representations in real images can be formulated as:
$\mathcal{L}_{hom}=\max\arg_{i\in I,k\in K^{\prime}(i)}\parallel z^{hom}_{k}-z^{hom}_{i}\parallel^{2}_{F},$ (2)
@@
Considering the potential issues of high feature coupling in the embedding space, we further exploit soft orthogonality to reduce information redundancy and dependencies between these homogeneous and heterogeneous features:
$\mathcal{L}_{ort}=\sum_{i\in I}\cos(z^{hom}_{i},z^{het}_{i}),$ (3)
@@
$\mathcal{L}_{tra}=\mathcal{L}_{hom}+\mathcal{L}_{het}+\lambda\mathcal{L}_{ort},$ (4)
@@
4.2 Fake Image Detection
To capture the homogeneous features as an additional predictive target, we employ an extended supervised contrastive learning to simultaneously encode real and fake images further for fake image detection. Specifically, we incorporate these homogeneous features into extended supervised contrastive learning and constrain their distance from positive and negative samples. This will motivate the detector to identify fake images based on distance. Note that the feature encoder is frozen during this stage.
Formulation. We now assume access a full dataset, $D=D_{r}\bigcup D_{f}$, where $D_{r}$ is used in the previous stage, $D_{f}$ is a dataset of fake images. Our architecture consists of a feature encoder, an auxiliary projection head for extended supervised contrastive learning, and another classification head for supervised classification(shown in Figure 2(b). With the auxiliary projection head, the real and fake feature embeddings are mapped to $z^{r}_{i}$ and $z^{f}_{i}$, respectively. To motivate the network to focus on the intra-class aggregation of real images more than the inter-class differences between real and fake images, we adopt the homogeneous features from the previous stage as extra positive instances for real anchors and additional negative instances for fake anchors. Formally, the extended supervised contrastive loss is as follows:
$\mathcal{L}_{ext}=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\frac{exp(z_{i}\cdot z_{p}/\tau)}{\sum_{k\in K(i)}exp(z_{i}\cdot z_{k}/\tau)},$ (5)
where $P(i)$ is the set of all positive samples for the anchor $i$, $|P(i)|$ is its cardinality. Note that for real anchor, the auxiliary homogeneous feature $z^{hom}\in P(i)$, and vice versa, $z^{hom}\in K(i)$. To achieve fake image detection, for a given sample and label, the classifier $D$ is optimized on the binary cross entropy loss:
@@ 4.2 Fake Image Detection
$\mathcal{L}_{d}=\mathcal{L}_{ext}+\gamma\mathcal{L}_{ce},$ (7)
@@
Figure 4: Ablation study on training data. All detectors trained on different data sources (ProGAN or LDM) or different numbers of classes of the ProGAN data source (20 classes, 12 classes, and 4 classes). The diversity of training data improves generalization ability.
@@
Table 5: Evaluation on commercial generative models. NTF shows 78.41% Acc on Midjourney, outperforming other baselines.
5 Experiments
5.1 Experiments Setup
Dataset. We use the dataset provided by Wang et al. (2020), which consists of 720K images for training and 4K images for validation. The fake images were generated by ProGANKarras et al. (2017), while an equal number of real images were sourced from the LSUNYu et al. (2015). We also conducted evaluation experiments on the self-built dataset, which covered GAN-based, DM-based, and multi-step fake images.
@@ 5.1 Experiments Setup
Baselines. For better evaluation, we compare with six baselines, including fake detection methods based on explicit artifacts and implicit artifacts. In experiments, we compare with the following six baselines: 1) XceptionNet Rossler et al. (2019) is widely employed as the baseline in the studies of DeepFake forensics; 2) Wang2020Wang et al. (2020) focuses on the artifacts exposed by CNN-generated images; 3) Grag2021Gragnaniello et al. (2021) uses spectral super-resolution to reconstruct visual cues for detection. 4) Ojha2023Ojha et al. (2023) uses a feature space not explicitly trained to distinguish real from fake images. 5) Self-BlendedShiohara and Yamasaki (2022) mixes image pairs with various masks to generate training data. 6) CADDMDong et al. (2023) focuses on local information so that the network ignores identity information leakage caused by irregular face changes. Except that Self-Blended and CADDM are trained on FaceForensics++ Rossler et al. (2019), the training set for the other comparison methods is consistent with ours. Note that Self-Blended and CADDM are limited to DeepFake datasets, and for the sake of fairness, they are not included in the test of non-face data.
Implementation Details. We use ResNet50 pre-trained on ImageNet as the feature encoder. All the projection heads contain two layers of MLPs with an output dimension of 128. For all datasets, we use a $224\times 224$ crop for both training and testing (random crop for training and center crop for testing). For extended supervised contrastive learning, we perform a random crop of the input image to 32 px. The optimizer is an SGD with a momentum of 0.9, an initial learning rate of 0.1, and an attenuation of 0.001. In the first stage, training is conducted for 200 epochs, followed by 10 epochs in the second stage. The empirical setting for $\lambda$, $\gamma$ is set at 0.1, 0.5.
@@
5.2 Effectiveness Evaluation
We explore the capability of NTF in detecting fake images from different generative models under the aforementioned experimental setup. Table 2-4 highlights its advanced performance through comparison with baselines.
@@ 5.2 Effectiveness Evaluation
Performance on GAN-based fake images. Experimental results in Table 2 demonstrate that NTF shows its capability in tackling unknown GAN-based fake images with the highest AP 99.91% and the lowest FPR 0.23% on average. In particular, NTF attains optimal detection performance on GAN-based fake images, which could be attributed to the similarity in the image generation principle of these models to that of ProGAN. That is precisely why certain baselines, such as XceptionNetRossler et al. (2019) and Grag2021Gragnaniello et al. (2021), have achieved high-performance generalization on GAN-based generative models. This indicates that existing fake image detection methods are adept at handling generalization scenarios within the same model family.
@@
Performance on DM-based fake images. As illustrated in Table 3, NTF exhibits superior cross-family generalization capability on the DMs with the highest AP 91.16% and ACC 82.67% on average. Specifically, NTF improves the AP and ACC by nearly 5.92% and 10.12% compared to the best baseline. The detection accuracy of NTF surpasses all baselines across four DMs. However, it remains comparable to Grag2021Gragnaniello et al. (2021) on LDM and to XeceptionNetRossler et al. (2019) on DDPM. Overall, NTF detects different DM-base generation models in a more balanced way, although it is slightly inferior to some baselines on certain models.
@@
Performance on multi-step fake images. As shown in Table 4, NTF detects fake images synthesized by six multi-step methods with the highest mAP of 97.32%. It is noteworthy that the method from Ojha2023Ojha et al. (2023) demonstrated an average ACC of 91.14%, slightly outperforming NTF. This can be attributed to their use of the ViT-L/14Dosovitskiy et al. (2020), a vision transformer variant pre-trained on CLIPRadford et al. (2021) for the backbone, aiding in effectively modeling details for real and fake image classification.
@@
Performance on Challenging Midjourney. To better evaluate the performance in tackling commercial generative models, we assess the detection capability of NTF on fake images generated by Midjourney in Table 5. Experimental results show that our proposed method NTF gives an accuracy more than 78% which significantly outperforms the baselines.
@@
In summary, NTF effectively detects GAN-based unknown generative models with a mAP 99.91%. It also demonstrates the capability to detect DM-based and multi-step generation methods with mAPs of 91.16% and 97.32% respectively, achieving an overall mAP of 96.18% across all datasets. Additionally, NTF shows a slightly higher FNR, which can be attributed to the inadequate coverage of real-world images in the training dataset.
5.3 Ablation Studies
@@
We evaluated the effect of architecture and training data on the generalization ability of NTF. It initially employed homogeneous, heterogeneous, and soft orthogonality losses with Eq (4) for learning natural trace representations and it trains with the ProGAN/LSUN dataset consisting of 20 classes.
@@
Effect of network architecture. We conducted experiments with different variants of NTF, exploring the following configurations: 1) without orthogonality loss, and 2) without heterogeneity and orthogonality loss. For each, We maintain ProGAN real/fake image data as training data. Figure 3 shows the performance of these variants on the same models. We find that the architectural design of NTF has an important role in generalization, improving performance on conditional generative models such as LDM, StableDM, and multi-step synthesis methods. This improvement is likely attributed to the implementation of soft orthogonality and heterogeneous constraints on upper bounds, which enables the NTF to learn homogeneous features capable of effectively separating traces of conditional generative models from natural traces. Moreover, this design also improves performance on unconditional generative models, particularly more significant in models like Diff-StyleGAN and PNDM.
@@
Effect of training data Next, we investigate how the source and diversity of the training data influence a detector’s generalization ability. Our approach involves altering either the data source or the number of classes in the training set. Specifically, we train multiple detectors by 1) employing a pre-trained LDM, substituting the ProGAN, and 2) utilizing a subset of the full ProGAN dataset, excluding real and fake images from certain LSUN classes, as shown in Figure 4. With access only to the LDM dataset, the model displays impressive generalization capabilities, even though the dataset consists of 200k reals from LAIONSchuhmann et al. (2022) and 200k fakes generated by LDM. As expected, diversifying the training set does enhance generalization to some extent, but the benefits diminish with increasing diversity. This suggests the potential existence of a real image dataset capable of extracting universally present natural traces.
5.4 Robustness Evaluation
@@
As shown in Figure 5, NTF is typically robust to blur, compression, and noise operations. The performance of Wang2020Wang et al. (2020) is more consistent under different levels of blur/compression/noise. This is reasonable, as it employs random compression and blur for data augmentation during training. Although Ojha2023Ojha et al. (2023) also employs data augmentation, it also performs in addition to being less robust in blur operations. In addition, both NTF and other baselines show significant performance degradation in image scaling operations. When the image scaling factor is 0.2, the image content becomes imperceptible, making robustness in such extreme scenarios less critical. Please refer to the technical appendix for more details.
5.5 Qualitative Analysis
To understand how the network generalizes to different synthesis methods, we visualize the model saliency map. We apply GradCAM++ Chattopadhay et al. (2018) to NTF on the self-built dataset to visualize where models are paying their attention to images, as shown in Figure 6. This demonstrates how the network captures artifacts from different generative models; for example, with the StyleGAN2-generated image of a dog, the network focuses on the inconsistency in shadow.
6 Conclusion
In this paper, we escape the trap of exploring subtle differences between real and fake for fake image detection. Motivated by the presence of common shared features in real images, We propose a novel framework named NTF, pre-trained by natural trace representation learning and extended supervised contrastive learning, to significantly improve the generalization ability of fake image detection. Extensive experiments on the self-built dataset demonstrate that our method exhibits state-of-the-art generalization capability for unknown generative models. Our research also offers a fresh perspective on fake image detection, focusing on exploring stable detectable features rather than those that continuously change, thereby paving the way for future studies in this field.
7 Technical Appendix
In the technical appendix, we present the details of the self-built dataset, evaluation of the test set from Wang et al. (2020), some visualizations, and additional experiments.

### v2

@@ A Diverse Generated Image Dataset
Formulation. We assume access to a large-scale dataset of real images $D_{r}$. Sample $x\in D_{r}$ is an augmented sample from an image $x_{r}\in\mathbb{R}^{H\times W\times 3}$. As shown in Figure 2(a), our architecture consists of a feature encoder followed by two projection heads that map real image embeddings into homogeneous and heterogeneous features.
Specifically, feature encoder, $E$, producing feature embedding $e=E(x_{r})$ from the inputs, then decouples the $e$ into homogeneous and heterogeneous features, $z^{hom}=f^{hom}(e)\in\mathbb{R}^{C}$ and $z^{het}=f^{het}(e)\in\mathbb{R}^{C}$ through two projection heads, $f^{hom}$ and $f^{het}$, respectively, where $C$ is the dimensionality of the features. Let $i\in I\equiv\{1...2N\}$ be the index of an arbitrary augmented sample, where there are a total of N samples in a batch, each with two random augmentations ($2N$). The self-supervised contrastive loss of heterogeneous (het.) feature representations in real images can be formulated as follows:
@@
where the index $i$ is called the anchor, the index $p$ is called the positive, $\tau$ is a temperature hyperparameter and $K(i)\equiv I\backslash\{i\}$, which contains all the augmented samples except $i$ in a batch. Furthermore, the loss of homogeneous (hom.) feature representations in real images can be formulated as:
$\mathcal{L}_{hom}=\arg\max_{i\in I,k\in K^{\prime}(i)}\parallel z^{hom}_{k}-z^{hom}_{i}\parallel^{2}_{F},$ (2)
@@
Considering the potential issues of high feature coupling in the embedding space, we further use soft orthogonality (orth.) to reduce information redundancy and dependencies between these homogeneous and heterogeneous features:
$\mathcal{L}_{orth}=\sum_{i\in I}\cos(z^{hom}_{i},z^{het}_{i}),$ (3)
@@
$\mathcal{L}_{tra}=\mathcal{L}_{hom}+\mathcal{L}_{het}+\lambda\mathcal{L}_{orth},$ (4)
@@
Transfer Learning for Fake Image Detection
To capture the natural traces as an additional target, we develop the SCL to simultaneously encode real and fake images further for fake image detection. Specifically, we incorporate these natural traces into SCL and constrain their distance from positive and negative samples. This will motivate the detector to identify fake images based on distance. Note that the feature encoder is frozen during this stage.
Formulation. We now assume access a full dataset, $D=D_{r}\bigcup D_{f}$, where $D_{r}$ is used in the previous stage, $D_{f}$ is a dataset of fake images. Our architecture consists of a feature encoder, an auxiliary projection head for SCL, and another classification head for supervised classification(shown in Figure 2(b). With the auxiliary projection head, the real and fake feature embeddings are mapped to $z^{r}_{i}$ and $z^{f}_{i}$, respectively. To motivate the network to focus on the intra-class aggregation of real images more than the inter-class differences between real and fake images, we adopt the homogeneous features $z^{hom}$ from the previous stage as extra positive instances for real anchors and additional negative instances for fake anchors. SCL mitigates the negative impact of fake image features on the classifier by assigning weights to the homogeneous features. Formally, the soft contrastive loss is as follows:
$\mathcal{L}_{scl}=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\sum_{j\in J}\frac{exp[z_{i}(z_{p}+y\cdot\eta\cdot z^{hom}_{j})/\tau]}{\sum_{k\in K(i)}exp(z_{i}\cdot z_{k}/\tau)},$ (5)
where $P(i)$ is the positive samples set for the anchor $i$, $|P(i)|$ is its cardinality, $K(i)$ is the negative set and $J$ is homogeneous set in a batch. Note that for real anchor, $y=1$, otherwise $y=-1$. $\eta$ is a balance factor. To achieve fake image detection, for a given sample and label, the classifier $D$ is optimized on the binary cross entropy loss:
@@
$\mathcal{L}_{d}=\mathcal{L}_{scl}+\gamma\mathcal{L}_{ce},$ (7)
@@
Figure 4: Ablation study on training data. All detectors trained on different data sources (ProGAN or LDM) or different numbers of classes of the ProGAN data source (20 classes, 12 classes, and 4 classes).
@@
Kolors/Acc(%) $\uparrow$ 48.26 60.24 57.69 78.12
Table 5: Evaluation on commercial generative models.
Experiments
Experiments Setup
Dataset. We use the dataset provided by (Wang et al. 2020), which consists of 720K images for training and 4K images for validation. The fake images were generated by ProGAN(Karras et al. 2017), while an equal number of real images were sourced from LSUN(Yu et al. 2015). We conduct evaluation experiments on the self-built dataset, which covered GAN-based, DM-based, and multi-step fake images.
@@
Baselines. We compare with the following six baselines, including fake detectors based on explicit and implicit artifacts: 1) Xception (Rossler et al. 2019) is widely employed as the baseline in the studies of DeepFake forensics; 2) Wang2020(Wang et al. 2020) focuses on the artifacts exposed by CNN-generated images; 3) Grag2021(Gragnaniello et al. 2021) uses spectral super-resolution to reconstruct visual cues for detection. 4) Ojha2023(Ojha 2023) uses a feature space not explicitly trained to distinguish real from fake images. 5) SBIs(Shiohara 2022) mixes image pairs with various masks to generate training data. 6) CADDM(Dong et al. 2023) focuses on local information so that the network ignores identity information leakage caused by irregular face changes. Except that SBIs and CADDM are trained on FaceForensics++ (Rossler et al. 2019), the training set for the other baselines is consistent with ours. Note that SBIs and CADDM are limited to DeepFake datasets, and for the sake of fairness, they are not included in the test of non-face data.
Implementation Details. We use ResNet50 pre-trained on ImageNet as the feature encoder. All the projection heads contain two layers of MLPs with an output dimension of 128. For all datasets, we use a $224\times 224$ crop for both training and testing (random crop for training and center crop for testing). For soft contrastive learning, we perform a random crop of the input image to 32 px. The optimizer is an SGD with a momentum of 0.9, an initial learning rate of 0.1, and an attenuation of 0.001. In the first stage, training is conducted for 200 epochs, followed by 10 epochs in the second stage. The empirical setting for $\lambda$, $\eta$, $\gamma$ is set at 0.1, 0.5, 0.5.
@@
Effectiveness Evaluation
We explore the capability of NTF to detect fake images from different generative models in the self-built dataset.
@@
Performance on GAN-based fake images. The results in Table 2 demonstrate the capability of NTF in tackling unknown GAN-based fake images with the highest AP 99.9% and the lowest FPR 0.2% on average. In particular, NTF attains optimal performance on GAN-based fake images, which could be attributed to the similarity in the image generation principle of these models to that of ProGAN. That is precisely why certain baselines, such as Xception and Grag2021, have achieved high-performance generalization on GAN-based generative models. This indicates that existing fake image detections are adept at handling generalization scenarios within the same model family.
@@
Performance on DM-based fake images. As illustrated in Table 3, NTF exhibits superior cross-family generalization capability on the DMs with the highest AP 91.2% and ACC 82.7% on average. Specifically, NTF improves the AP and ACC by nearly 6% and 19.2% compared to the best baseline. The detection accuracy of NTF surpasses all baselines across four DMs. However, it remains comparable to Grag2021 on LDM and to XeceptionNet on DDPM. Overall, NTF detects different DM-base generation models in a more balanced way, although it is slightly inferior to some baselines on certain models.
@@
Performance on multi-step fake images. As shown in Table 4, NTF detects fake images synthesized by six multi-step methods with the highest mAP of 97.3%. It is noteworthy that the method from Ojha2023 demonstrated an average ACC of 91.1%, slightly outperforming NTF. This can be attributed to the use of ViT-L/14(Dosovitskiy et al. 2020), a vision transformer variant pre-trained on CLIP(Radford et al. 2021) for the backbone, aiding in effectively modeling details for real and fake image classification.
@@
Performance on commercial generative models. To better evaluate the performance in tackling commercial generative models, we assess the detection capability of NTF on fake images generated by Midjourney and Kolors in Table 5. Experimental results show that our proposed method NTF gives an accuracy more than 78.4% which significantly outperforms the baselines.
@@
In summary, NTF effectively detects GAN-based unknown generative models with a mAP 99.9%. It also demonstrates the capability to detect DM-based and multi-step generation methods with mAPs of 91.2% and 97.3% respectively, achieving an overall mAP of 96.2% across all datasets. Additionally, NTF shows a slightly higher FNR, which can be attributed to the inadequate coverage of real-world images in the training dataset.
Ablation Studies
@@
We evaluated the effect of architecture and training data on the generalization ability of NTF. It initially employed homogeneous, heterogeneous, and soft orthogonality losses with Eq (4) for learning natural trace representations and it trains with ProGAN/LSUN dataset consisting of 20 classes.
@@
Effect of network architecture. We conducted experiments with different variants of NTF, exploring the following configurations: 1) without ort. loss, 2) without het. and orth. loss and 3) without SCL. For each, We maintain ProGAN real/fake image data as training data. Figure 3 shows the performance of these variants on the same models. We find that the architectural design of NTF has an important role in generalization, improving performance on conditional generative models such as LDM, SDM, and multi-step synthesis methods. This improvement is likely attributed to the implementation of soft orthogonality and heterogeneous loss constraints on upper bounds, which enables the NTF to learn homogeneous features capable of effectively separating traces of conditional generative models from natural traces. Moreover, this design also improves performance on unconditional generative models, particularly more significant in models like Diff-StyleGAN and PNDM.
@@
Effect of training data Next, we investigate how the source and diversity of the training data influence a detector’s generalization ability. Our approach involves altering either the data source or the number of classes in the training set. Specifically, we train multiple detectors by 1) employing a pre-trained LDM, substituting the ProGAN, and 2) utilizing a subset of the full ProGAN dataset, excluding real and fake images from certain LSUN classes, as shown in Figure 4. With access only to the LDM dataset, the model displays impressive generalization capabilities, even though the dataset consists of 200k reals from LAION(Schuhmann et al. 2022) and 200k fakes generated by LDM. As expected, diversifying the training set does enhance generalization to some extent, but the benefits diminish with increasing diversity. This suggests the potential existence of a real image dataset capable of extracting universally present natural traces.
Robustness Evaluation
@@
As shown in Figure 5, NTF is typically robust to blur, compression, and noise operations. The performance of Wang2020 is more consistent under different levels of blur/compression/noise. This is reasonable, as it employs random compression and blur for data augmentation during training. Although Ojha2023 also employs data augmentation, it also performs in addition to being less robust in blur operations. In addition, both NTF and other baselines show significant performance degradation in image scaling operations. When the image scaling factor is 0.2, the image content becomes imperceptible, making robustness in such extreme scenarios less critical. Please refer to the technical appendix for more details.
Qualitative Analysis
To understand how the network generalizes to different synthesis methods, we visualize the model saliency map. We apply GradCAM++ (Chattopadhay et al. 2018) to NTF on the self-built dataset to visualize where models are paying their attention to images, as shown in Figure 6. This demonstrates how the network captures artifacts from different generative models; for example, with the dog generated by StyleGAN2, the network focuses on the inconsistency in shadow.
Conclusion
In this paper, we escape the trap of exploring subtle differences between real and fake for fake image detection. Motivated by the presence of common shared features in real images, We propose a novel framework named NTF, pre-trained by natural trace representation learning and soft contrastive learning, to significantly improve the generalization ability of fake image detection. Extensive experiments on the self-built dataset demonstrate that our method exhibits state-of-the-art generalization capability for unknown generative models. Our research also offers a fresh perspective on fake image detection, focusing on exploring stable detectable features rather than those that continuously change, thereby paving the way for future studies in this field.
Acknowledgements
This research was supported in part by the National Key Research and Development Program of China under No.2021YFB3100700, the National Natural Science Foundation of China (NSFC) under Grants No. 62202340, 62372334, the CCF-NSFOCUS ‘Kunpeng’ Research Fund under No. CCF-NSFOCUS 2023005, the Open Foundation of Henan Key Laboratory of Cyberspace Situation Awareness under No. HNTS2022004, the Fundamental Research Funds for the Central Universities under No. 2042023kf0121.